OpenStudy (anonymous):

can anyone PLEASE HELP ME I have 6 more problems

OpenStudy (anonymous):

12 + 7( 9 ÷ 3) + 4

OpenStudy (mertsj):

12+7(3)+4= 12+21+4=37

OpenStudy (anonymous):

the other is simplify: -2b(3a – 2b) + 4(3ab – 2b2)

sam (.sam.):

-2b(3a-2b)+4(3ab-2b^(2)) -6ab+4b^(2)+4(3ab-2b^(2)) -6ab+4b^(2)+12ab-8b^(2) 6ab-4b^(2)

OpenStudy (anonymous):

6(4 + 2x) – 3( -2 – 4x)

OpenStudy (mertsj):

24+12x+6+12x=30+24x

OpenStudy (anonymous):

Simplify: 2a – {6b – 4[ a – (b – 3a)]}

OpenStudy (mertsj):

2a-{6b-4[a-b+3a]} 2a-{6b-4a+4b-12a} 2a-{10b-16a} 2a-10b+16a 18a-10b